OpenStudy (anonymous):

Help pleaseee<3 What is the surface area of a conical grain storage tank that has a height of 24 meters and a diameter of 16 meters? round the answer to the nearest square meter. 837 square meters 848 square meters 1,991 square meters 1,923 square meters

OpenStudy (tkhunny):

\(2\pi r^{2} + 2\pi r h\) Go!! Lids: \(2\pi r^{2}\) Sides: \(2\pi r h\)

OpenStudy (anonymous):

well he gave you formula what else'd you need?

OpenStudy (anonymous):

Formula: http://www.mathwarehouse.com/solid-geometry/cone/images/surface-area-of-cone/total-surface-formula-cone.jpg \(\Large S_a=πrh_s+πr^2=πr\sqrt{r^2+h^2}+πr^2 \)